# Deferred Poisoning: Making the Model More Vulnerable via Hessian Singularization

### v2

@@ IV-E Robustness analysis
TABLE V: The accuracy and robustness(%) of poisoned models over CIFAR10 after data augmentation.
Method Robustness VGG16 ResNet18 ResNet50
@@
IV-F Different poisoning percentages
@@
TABLE VI: The accuracy and robustness(%) of VGG16 under different poisoning percentages in CIFAR10.
Robustness Percentage 0% 20% 40% 60% 80% 100%
@@
Table VI displays the performance of the VGG16 model under various percentages of deferred poisoning attacks on the CIFAR10 dataset. As the poisoning ratio increases, the model becomes increasingly vulnerable. When the poisoning ratio reaches $40\%$, the deferred poisoning attack has already demonstrated its offensiveness. Even at a ratio of $60\%$, the model suffers significant harm, which further illustrates the severe threat that deferred poisoning attacks pose to machine learning, poisoning only a portion of the training data can pose a serious risk to the model.
IV-G Hessian-vector product performance
In Section III-D, we introduce the Hessian-vector product (HVP) to calculate the relaxed solution of the Hessian singularization term more quickly. With the new strategy, it now takes about 240 minutes to poison the entire CIFAR10, and about 30 hours to Tiny-ImageNet for large-scale scenarios, which is comparable to the competitive methods in our paper as shown in Table VII.
TABLE VII: The comparison of the robustness(%) compared with the origin method and HVP.
@@ IV-G Hessian-vector product performance
For example, as shown in the 1st-3rd row of the 1st column of Table VII, when training CIFAR10 with VGG16 after being attacked by the original DPA, the model’s robustness under FGSM attack drops from 1.86% to 0.54%. In contrast, when training with the dataset under HVP rapid attack, the robustness is 0.58%, which is close to the original attack method. This indicates that HVP can accelerate the poisoning process without compromising the effectiveness of the DPA attack.
IV-H Adversarial training
Adversarial training has emerged as a pivotal technique for enhancing the robustness of models, particularly when it comes to safeguarding against adversarial attacks. For example, Madry et al. [18] train the model to minimize loss function under adversarial perturbations by using Projected Gradient Descent (PGD) as an adversary, thereby enhancing the model’s robustness against a broad range of adversarial attacks. Foret et al. [36] enhance model’s generalization by concurrently reducing both its loss value and the sharpness of the loss landscape.
TABLE VIII: The comparison of the accuracy and robustness(%) of different kinds of adversarial training method against DPA.
VGG16-CIFAR10 ResNet18-CIFAR10 ResNet50-CIFAR10
Clean 0.89 1.86 0.94 1.19 0.81 1.62 1.01 1.24 0.78 2.84 1.48 1.21
No defense 0.75 0.58 0.50 0.32 0.76 0.79 0.57 0.39 0.74 1.38 0.62 0.45
SAM 0.76 0.57 0.50 0.30 0.79 0.84 0.57 0.41 0.78 0.98 0.57 0.39
AT 0.78 0.71 0.77 0.94 0.79 1.07 0.78 0.79 0.78 1.51 0.86 0.87
Ours 0.80 1.72 1.25 1.52 0.71 1.93 1.31 1.54 0.68 2.51 1.54 1.83
Table VIII demonstrates the effectiveness of DPA attacks against various adversarial training methods, where AT denotes adversarial training using PGD [18]. As shown in the 4th row of the 6th column, traditional adversarial training methods are not entirely effective in defending against our attacks. For instance, when training CIFAR10 with VGG16 and under the FGSM attack, the robustness of the clean model is 1.86%, whereas the robustness of the model poisoned after adversarial training is only 0.71%.
@@ IV-H Adversarial training
The results also show that SAM is ineffective against our attack method. For example, the model robustness of VGG16-CIFAR10, with and without SAM adversarial training, is nearly the same, 0.57% (the 4th row of the 3rd column) and 0.58% (the 5th row of the 3rd column), respectively. SAM is ineffective possibly because it focuses on decreasing the curvature of the loss function with respect to model parameters. In contrast, our method aims to increase the curvature of the loss function concerning the input examples. They are two different dimensionalities of the loss function. Therefore, SAM cannot effectively counter our attack method.
@@
Consequently, we propose a new training paradigm that directly decreasing the curvature of the loss function with respect to the input(namely, Ours in Table VIII). For example, in the 2nd row and the 5th row of the table, the robustness increases from 0.58% to 1.72%. This finding warns researchers that simply adopting commonly used adversarial training strategies is not enough to ensure a trustworthy model. Considering the curvature information maliciously used by our DPA is also crucial. In summary, our work focuses on revealing a new attack. Like most pioneering works (e.g., FGSM, PGD) in adversarial machine learning, we present a novel attack method, and a defense method naturally follows to design a more robust model.
V Conclusion
In this paper, we introduce the Deferred Poisoning Attack (DPA), a stealthy threat in artificial intelligence that performs normally during training and validation. DPA targets the Hessian matrix, inducing a rapid convergence to a steep local optimum on poisoned data, thereby diminishing the model’s robustness. We demonstrate DPA’s effectiveness and stealth through comparative adversarial analyses with traditional attacks. Notably, DPA demands lower costs and exhibits transferability. We also establish its resilience against common augmentations and pretrained defenses, highlighting its universal risk underscores a novel challenge for AI security.

### v3

@@ Introduction
Table 5: The accuracy and robustness(%) of poisoned ResNet18 over CIFAR10 after data augmentation.
Adversarial training
Table 10 demonstrates the effectiveness of DPA attacks against various adversarial training methods, where AT denotes adversarial training using PGD and SAM, Smoothing, TRADES are proposed by (foret2021sharpnessaware; cohen2019certified; zhang2019theoretically) respectively. As shown in the table, traditional adversarial training methods are not entirely effective in defending against our attacks. Consequently, we propose a new training paradigm that directly decreasing the curvature of the loss function with respect to the input(namely, Ours).
Method ACC $\hat{\rho}_{F}\uparrow$ $\hat{\rho}_{P}\uparrow$ $\hat{\rho}_{D}\uparrow$
Clean 0.81 1.62 1.01 1.24
No defense 0.76 0.79 0.57 0.39
SAM 0.79 0.84 0.57 0.41
AT 0.79 1.07 0.78 0.79
Smoothing 0.75 1.11 0.80 0.92
TRADES 0.78 1.19 0.82 0.85
Ours 0.71 1.93 1.31 1.54
Table 6: The comparison of the accuracy and robustness(%) of different adversarial training methods against DPA.
This finding warns researchers that simply adopting commonly used adversarial training strategies is not enough to ensure a trustworthy model. Considering the curvature information maliciously used by our DPA is also crucial. Like most pioneering works in adversarial machine learning, we present a novel attack method, and a defense method naturally follows to design a more robust model.
Conclusion
In this paper, we introduce the DPA, a stealthy threat in AI that performs normally during training and validation. DPA targets the Hessian matrix, inducing a rapid convergence to a steep local optimum on poisoned data, thereby diminishing the model’s robustness. We demonstrate DPA’s performance through comparative adversarial analyses with traditional attacks. Notably, DPA demands lower costs and exhibits transferability. We also establish its resilience against common augmentations and pretrained defenses, highlighting its universal risk underscores a novel challenge for AI security.
Acknowledgments
This research was partially supported by the National Natural Science Foundation of China (Grant Nos. 62202009 and 62571284), the Macau Science and Technology Development Fund (Grant Nos. 0040/2023/ITP1 and 0004/2023/RIB1), and the Basic and Applied Basic Research Foundation of Guangdong Province (Grant No. 2024A1515011755).
Appendix A Appendix
Appendix B Complexity analysis
Here is a specific analysis of computational complexity:
Time complexity
  • Stage 1. Train Model: Update model parameters: $O(MP)$ (shah2022time)
  • Stage 2. Update perturbation
    • Step 1. Perform a forward pass to obtain the Loss: $O(MP)$
    • Step 4. Compute $\text{tr}(H^{T}H)$ or $\|Hv\|_{F}$: $O(P)$
    • Step 5. Backpropagate and update $\epsilon$: $O(P)$
Space complexity
Appendix C Experiment
Experiment details
Table 7 presents the hyperparameter settings relevant to the generation of DPA perturbations in our experiments.
Parameter Value Explanation
seed 0 Random seed.
lr 0.01 Learning rate of model training.
num_epochs 20 Iterations of model training.
batch_size 64 The batch size of training data.
optimizer SGD The optimizer of model training.
optimizer_epsilon Adam The optimizer of perturbation updating.
lr_epsilon 0.001 Learning rate of perturbation updating.
num_epochs_epsilon 20 Iterations of perturbation updating.
Table 7: Hyperparameter Settings for Generating Perturbations.
Figure 6: Image after adding different random noise
Method VGG16 ResNet18 ResNet50
@@ Appendix C Experiment
Table 8: The accuracy and robustness(%) of poisoned models over CIFAR10 after data augmentation.
Visualization of poisoned samples
Visualization of nature noise
We experimented with natural noise from various distributions to simulate real-world scenarios as show in Fig. 6:
Robustness
Table 8 completely displays the performance of deferred poisoning attacks when facing various data augmentation techniques with $\epsilon$ set to $3/255$. It is evident that our method can still render models vulnerable under most data augmentation methods as well as with pre-trained models, indicating that our approach possesses strong robustness.
Different poisoning percentages
@@
Robustness 0% 20% 40% 60% 80% 100%
@@
Table 9: The accuracy and robustness(%) of VGG16 under different poisoning percentages in CIFAR10.
Table 9 displays the performance of the VGG16 model under various percentages of deferred poisoning attacks on the CIFAR10 dataset. As the poisoning ratio increases, the model becomes increasingly vulnerable. When the poisoning ratio reaches $40\%$, the deferred poisoning attack has already demonstrated its offensiveness. Even at a ratio of $60\%$, the model suffers significant harm, which further illustrates the severe threat that deferred poisoning attacks pose to machine learning, poisoning only a portion of the training data can pose a serious risk to the model.
Adversarial training
Table 10 completely demonstrates the effectiveness of DPA attacks against various adversarial training methods, where AT denotes adversarial training using PGD (PGD). As shown in the 6th row, traditional adversarial training methods are not entirely effective in defending against our attacks. For instance, when training CIFAR10 with VGG16 and under the FGSM attack, the robustness of the clean model is 1.86%, whereas the robustness of the model poisoned after adversarial training is only 0.71%.
VGG16-CIFAR10 ResNet18-CIFAR10 ResNet50-CIFAR10
Clean 0.89 1.86 0.94 1.19 0.81 1.62 1.01 1.24 0.78 2.84 1.48 1.21
No defense 0.75 0.58 0.50 0.32 0.76 0.79 0.57 0.39 0.74 1.38 0.62 0.45
SAM 0.76 0.57 0.50 0.30 0.79 0.84 0.57 0.41 0.78 0.98 0.57 0.39
AT 0.78 0.71 0.77 0.94 0.79 1.07 0.78 0.79 0.78 1.51 0.86 0.87
Ours 0.80 1.72 1.25 1.52 0.71 1.93 1.31 1.54 0.68 2.51 1.54 1.83
Table 10: The comparison of the accuracy and robustness(%) of different kinds of adversarial training method against DPA.
@@
Table 11: The comparison of the robustness(%) compared with the origin method and HVP.
@@
The results also show that SAM is ineffective against our attack method. For example, the model robustness of VGG16-CIFAR10, with and without SAM adversarial training, is nearly the same, 0.57% (the 5th row of the 3rd column) and 0.58% (the 4th row of the 3rd column), respectively. SAM is ineffective possibly because it focuses on decreasing the curvature of the loss function with respect to model parameters. In contrast, our method aims to increase the curvature of the loss function concerning the input examples. They are two different dimensionalities of the loss function. Therefore, SAM cannot effectively counter our attack method.
@@
Consequently, we propose a new training paradigm that directly decreasing the curvature of the loss function with respect to the input(namely, Ours in Table 10). For example, in the 4th row and the 7th row of the table, the robustness increases from 0.58% to 1.72%. This finding warns researchers that simply adopting commonly used adversarial training strategies is not enough to ensure a trustworthy model. Considering the curvature information maliciously used by our DPA is also crucial. In summary, our work focuses on revealing a new attack. Like most pioneering works (e.g., FGSM, PGD) in adversarial machine learning, we present a novel attack method, and a defense method naturally follows to design a more robust model.
Hessian-vector product performance
In the paper, we introduce the Hessian-vector product (HVP) to calculate the relaxed solution of the Hessian singularization term more quickly. With the new strategy, it now takes about 240 minutes to poison the entire CIFAR10, and about 30 hours to Tiny-ImageNet for large-scale scenarios, which is comparable to the competitive methods in our paper as shown in Table 11.
For example, as shown in the 1st-3rd row of the 1st column of Table 11, when training CIFAR10 with VGG16 after being attacked by the original DPA, the model’s robustness under FGSM attack drops from 1.86% to 0.54%. In contrast, when training with the dataset under HVP rapid attack, the robustness is 0.58%, which is close to the original attack method. This indicates that HVP can accelerate the poisoning process without compromising the effectiveness of the DPA attack.